Compute the relative rate of diffusion of nitric oxide (NO) to ammonia (NH3)
1
Expert's answer
2014-12-22T09:11:34-0500
Therelative rate of diffusion will be sqrt (M(NH3)/M(NO)) = sqrt (17/30) = 0.753